Complete the following one factor ANOVA table.

ANOVA

Source of Variation

SS

df

MS

F-test

Between Groups

3

Within Groups

405

Total

888

31

What is the F-test statistic?

Group of answer choices

11.13

4.45

20.51

2.06
Answer #1

From given ANOVA table we get,

Between groups SS = Total SS - Within grouos SS = 888 - 405

=> Between groups SS = 483

Within Groups df = Total df - Between groups df = 31 - 3 = 28

Between groups MS = 483 / 3 = 161

Within Groups MS = 405 / 28 = 14.46

F-test statistic is,

F = Between groups MS / Within groups MS = 161 / 14.46 =11.13

=> F-test statistic = 11.13